OpenStudy (dannyrod2000):

A(b)=10xb+7/2 the function A(b) relates the area of a trapezoid with a given height of 10 and one base length of 7 with the lengths of its other base which equation below represents the inverse function B(a) which takes the trapezoid's area as input and returns as output the length of the other base? A. a/7-5 B. a/5+7 C. a/5-7 D. a/7+5

OpenStudy (dannyrod2000):

@IAmSinged

OpenStudy (anonymous):

I think you messed up the problem and the answers, as I an getting completely different answers when I'm using PEMDAS ordering.

OpenStudy (anonymous):

Take a screenshot if you can, and I'll work from there.

OpenStudy (dannyrod2000):

um i had trouble putting pictures

OpenStudy (anonymous):

Printscreen=>post in paint=>save it =>load to imgur

OpenStudy (dannyrod2000):

i could instead right all the information down

OpenStudy (anonymous):

Alright, I worked out what the problem initially was. A(b) = 10 * (b + 7)/2 A(b) = 5 *(b + 7) A(b) = 5b + 35 5b = A(b) - 35 b = A(b)/5 - 7 Switch variables: B(a) = a/5 - 7 C. In the future, please keep in mind the syntax of a problem. If you change the syntax to something more comfortable-looking to you, you can ruin the PEMDAS ordering of the problem, making the problem completely different.

OpenStudy (dannyrod2000):

thanks it really helped @IAmSinged